tan(α+β) =3 and tan (α-β)=2 then value of sin 2α is

Grade:11

1 Answers

vishesh jain
6 years ago
Tan2A = tan( (A+B) + (A-B) ) = 3+2/1-3*2=-1 , 1+ tan^2 2A= sec^ 2A , cos2A= 1/(2)^1/2 ,Sin2A= tan2A*cos2A= -1/(2)^1/2